Find whether the following series are convergent or not


ii. ∞Σn=1 (√(n^2+3) - √(n^2-3)/ √n

1
Expert's answer
2022-02-07T15:58:15-0500
"\displaystyle\sum_{n=2}^{\infin}\dfrac{\sqrt{n^2+3}-\sqrt{n^2-3}}{n}"

"=\displaystyle\sum_{n=2}^{\infin}\dfrac{n^2+3-n^2+3}{n(\sqrt{n^2+3}+\sqrt{n^2-3})}"


"=\displaystyle\sum_{n=2}^{\infin}\dfrac{6}{n(\sqrt{n^2+3}+\sqrt{n^2-3})}"

Use Limit Comparison Test


"\lim\limits_{n\to\infin}\dfrac{a_n}{b_n}=\lim\limits_{n\to\infin}\dfrac{\dfrac{6}{n(\sqrt{n^2+3}+\sqrt{n^2-3})}}{\dfrac{1}{n^2}}=6,"

The "p" -series "\displaystyle\sum_{n=2}^{\infin}\dfrac{1}{n^2}" converges since "p=2>1."

Therefore the series "\displaystyle\sum_{n=2}^{\infin}\dfrac{\sqrt{n^2+3}-\sqrt{n^2-3}}{n}" is convergent by Limit Comparison Test.